# Homological stratification and descent

### v1

@@ 8. Comparison between h-stratification and tt-stratification
A theory of stratification for tt-categories is developed in [BHS23b] under the assumption that the Balmer spectrum is weakly noetherian. Our next goal is to relate this notion of stratification (based on the Balmer spectrum and the Balmer–Favi support) with the homological stratification developed in this paper (based on the homological spectrum and the homological support).
@@
We will say that a rigidly-compactly generated tt-category whose spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is tt-stratified if it is stratified in the sense of [BHS23b].
@@
$(a)\Rightarrow(b)$: By [BHS23a, Theorem 4.7] if $\mathscr{T}$ is tt-stratified, then the Nerves of Steel conjecture holds for $\mathscr{T}$, so it remains to show that $\mathscr{T}$ is h-stratified. It was proved in [BHS23a, Lemma 3.7] that $\operatorname{Supp}(E_{\mathscr{B}})=\{\pi(\mathscr{B})\}$ for any $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$. Hence $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle g_{\pi(\mathscr{B})}\rangle$ when $\mathscr{T}$ is tt-stratified. Therefore,
@@
for any $t\in\mathscr{T}$. This coincides with $\operatorname{Locid}\langle E_{\mathscr{B}}\mid\mathscr{B}\in\operatorname{Supp}^{h}(t)\rangle$ since [BHS23a, Theorem 4.7] establishes that $\pi^{-1}(\operatorname{Supp}(t))=\operatorname{Supp}^{h}(t)$ when $\mathscr{T}$ is tt-stratified. This establishes that $\mathscr{T}$ is h-stratified by 3.1.
$(b)\Rightarrow(a)$: If $\mathscr{T}$ is h-stratified then the h-detection property holds. Thus, 5.6 implies that $\pi(\operatorname{Supp}^{h}(t))=\operatorname{Supp}(t))$ for each $t\in\mathscr{T}$. It follows that the diagram
[…]
@@
For each $i\in I$, the tt-category $\mathscr{S}_{i}$ is tt-stratified by assumption, hence h-stratified by 8.6. Descent for h-stratification as established in 7.21 then implies that $\mathscr{T}$ is h-stratified. Therefore, 8.6 implies that $\mathscr{T}$ is tt-stratified, due to our assumption that $\mathscr{T}$ satisfies Nerves of Steel. ∎
@@
The equivalence between $(a)$ and $(b)$ is 8.7. 7.11 gives $(b)\Rightarrow(c)$, while the reverse implication is 7.21. ∎
@@ 9. Descent for the Nerves of Steel conjecture
where the vertical arrows are the finite localizations and the bottom arrow is the induced functor. This diagram satisfies the Beck–Chevalley condition that $h^{*}f_{*}\simeq g_{*}k^{*}$; cf. [San22, Lemma 5.5]. Indeed, this may be checked after applying the conservative $h_{*}$. For any $s\in\mathscr{S}$, we have $h_{*}h^{*}f_{*}(s)\simeq{f_{Y}\otimes f_{*}(s)}\simeq{f_{*}(f^{*}(f_{Y})\otimes s)}\simeq f_{*}(f_{\varphi^{-1}(Y)}\otimes s)\simeq f_{*}(k_{*}k^{*}(\mathbb{1})\otimes s)\simeq f_{*}k_{*}k^{*}(s)\simeq h_{*}g_{*}k^{*}(s)$ where we have invoked [BS17, Proposition 5.11]. With this in hand, consider a compact object $x\in\mathscr{S}(\varphi^{-1}(V))^{c}$. Then $x\oplus\Sigma x=k^{*}(c)$ for some $c\in\mathscr{S}^{c}$ and it suffices to show that $g_{*}k^{*}(c)$ is compact. This follows from our hypothesis, since $g_{*}k^{*}(c)\simeq h^{*}f_{*}(c)$. ∎
@@
[…]
Recall from 7.8 that $(f_{i}^{*})$ being jointly nil-conservative implies that $\varphi$ is also surjective. If $\varphi$ is in addition injective as assumed in $(a)$, then it is bijective. Since $\sqcup\pi_{\mathscr{S}_{i}}$ is bijective, $\varphi^{h}$ is injective and hence bijective. Therefore $\pi_{\mathscr{T}}$ is bijective.
Now assume either condition $(b)$ or $(c)$. By 6.14, the Nerves of Steel conjecture holds also for any finite localization of $\mathscr{S}_{i}$. Moreover, the property of being strongly closed is local in the target by 9.1, while the property of being nil-conservative is local in the target by 7.8[$(c)\Leftrightarrow(d)$] and 6.15. Therefore, we can reduce to the case when $\mathscr{T}$ is local and check that the fiber of $\pi_{\mathscr{T}}\colon\operatorname{Spc}^{h}(\mathscr{T}^{c})\to\operatorname{Spc}(\mathscr{T}^{c})$ over the unique closed point $\mathfrak{m}\in\operatorname{Spc}(\mathscr{T}^{c})$ is a singleton. Since $\varphi$ is surjective, we may choose $i\in I$ and a closed point $\mathscr{M}\in\operatorname{Spc}(\mathscr{S}_{i}^{c})$ lying over $\mathfrak{m}$. For the remainder of this proof, we set $\mathscr{S}=\mathscr{S}_{i}$ and simplify the notation accordingly.
@@
and hence $f^{*}(E_{\mathscr{B}})\otimes z=0$ because homological support detects weak rings (1.8). Therefore, $E_{\mathscr{B}}\otimes f_{*}(z)=0$ by the projection formula. That is $\mathscr{B}\not\in\operatorname{Supp}^{h}(f_{*}(z))$. By the hypothesis that $f$ is strongly closed, $f_{*}(z)$ is compact and hence $\operatorname{Supp}^{h}(f_{*}(z))=\pi_{\mathscr{T}}^{-1}(\operatorname{supp}(f_{*}(z)))$ by 5.2. Therefore we have $\mathscr{B}\not\in\pi_{\mathscr{T}}^{-1}(\operatorname{supp}(f_{*}(z)))$ so $\mathfrak{m}=\pi_{\mathscr{T}}(\mathscr{B})\not\in\operatorname{supp}(f_{*}(z))=\varphi(\operatorname{supp}(z))$ where the last equality is by [BCHS23, Theorem 13.13] since $z$ is a compact weak ring. But $\varphi(\operatorname{supp}(z))$ contains $\varphi(\{\mathscr{M}\})=\mathfrak{m}$ yielding the desired contradiction. ∎
@@
The results of 9.2 are not entirely satisfying; variations on the theme are possible and we expect a more complete statement to exist. Nevertheless, using 9.2$(a)$, we can provide a variant of the argument used in [Bal20b] to deduce the Nerves of Steel conjecture from nilpotence-type theorems in several examples of interest:
@@
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category. Suppose there exists a tt-field $f_{\mathscr{P}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{P}}$ detecting $\mathscr{P}$ for each prime $\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})$. The following are equivalent:
@@
[…]
@@
[…]
we see that $(a)$ implies that the map $\varphi^{h}\colon\bigsqcup_{\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})}\operatorname{Spc}^{h}(\mathscr{F}_{\mathscr{P}}^{c})\to\operatorname{Spc}^{h}(\mathscr{T}^{c})$ is surjective. This is equivalent to the family being nil-conservative by [BCHS24, Theorem 1.9]. Thus $(a)\Rightarrow(c)$. Finally, the implication $(c)\Rightarrow(a)$ follows from 9.2(a). ∎
@@
The next result establishes an equivalent formulation for the Nerves of Steel conjecture entirely in terms of the tensor triangular support. It strengthens previous work [BHS23a, Proposition 3.13] and could also be used to give an alternative deduction of 8.6.
@@
Assume first that the Nerves of Steel conjecture holds for $\mathscr{T}$. Using 5.10 twice as well as the tensor product formula for homological support (1.1), we compute
@@
Conversely, suppose that the tensor product formula holds for all weak rings. The argument in the proof of [BHS23a, Proposition 3.13] then implies the Nerves of Steel conjecture for $\mathscr{T}$. Note that the result there assumed the full tensor product formula, but it is only used for the weak rings $E_{\mathscr{B}}$. ∎
@@ 10. Applications to tt-stratification
We now state our main applications. We start with three descent results which extend those of [BCHS24, Section 17] and [BCH+24b, Section 12] to a family of functors.
@@
By 8.6 each $\mathscr{S}_{i}$ is h-stratified and satisfies the Nerves of Steel conjecture. The result then follows from 9.2 and 8.8 since $(f_{i}^{*})_{i\in I}$ is jointly nil-conservative by 7.8. ∎
@@
In the situation of 10.1$(c)$, if the indexing set $I$ is finite then $\operatorname{Spc}(\mathscr{T}^{c})$ is automatically noetherian and hence the hypothesis that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is superfluous.
@@
Note that the base-change functor $A\otimes-\colon\mathscr{C}\to\operatorname{Mod}_{\mathscr{C}}(A)$ induces a surjection $\operatorname{Spc}(\operatorname{Mod}_{\mathscr{C}}(A)^{c})\to\operatorname{Spc}(\mathscr{C}^{c})$, so $\operatorname{Spc}(\mathscr{C}^{c})$ is noetherian. Then we invoke 10.1$(c)$. ∎
@@
The previous corollary strengthens the results of [BCH+24b, Section 12] by removing the hypothesis that the descendable algebra $A$ be separable of finite tt-degree.
@@
[…]
in which the three indicated vertical maps are surjections. The top and middle horizontal maps are also surjections by 7.8. The bottom map is a bijection by assumption, hence $\rho_{\mathscr{T}}$ is surjective as well. Since the tt-categories $\mathscr{S}_{i}$ are cohomologically stratified, the left vertical composite is a bijection, which then implies that all maps in this diagram are bijections. The result thus follows from 8.8. ∎

### v2

@@ 8. Comparison between h-stratification and tt-stratification
A theory of stratification for tt-categories is developed in [BHS23d] under the assumption that the Balmer spectrum is weakly noetherian. Our next goal is to relate this notion of stratification (based on the Balmer spectrum and the Balmer–Favi support) with the homological stratification developed in this paper (based on the homological spectrum and the homological support).
@@
We will say that a rigidly-compactly generated tt-category whose spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is tt-stratified if it is stratified in the sense of [BHS23d].
@@
$(a)\Rightarrow(b)$: By [BHS23c, Theorem 4.7] if $\mathscr{T}$ is tt-stratified, then the Nerves of Steel conjecture holds for $\mathscr{T}$, so it remains to show that $\mathscr{T}$ is h-stratified. It was proved in [BHS23c, Lemma 3.7] that $\operatorname{Supp}(E_{\mathscr{B}})=\{\pi(\mathscr{B})\}$ for any $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{T}^{c})$. Hence $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle g_{\pi(\mathscr{B})}\rangle$ when $\mathscr{T}$ is tt-stratified. Therefore,
@@
for any $t\in\mathscr{T}$. This coincides with $\operatorname{Locid}\langle E_{\mathscr{B}}\mid\mathscr{B}\in\operatorname{Supp}^{h}(t)\rangle$ since [BHS23c, Theorem 4.7] establishes that $\pi^{-1}(\operatorname{Supp}(t))=\operatorname{Supp}^{h}(t)$ when $\mathscr{T}$ is tt-stratified. This establishes that $\mathscr{T}$ is h-stratified by 3.1.
$(b)\Rightarrow(a)$: If $\mathscr{T}$ is h-stratified then the h-detection property holds. Thus, 5.7 implies that $\pi(\operatorname{Supp}^{h}(t))=\operatorname{Supp}(t))$ for each $t\in\mathscr{T}$. It follows that the diagram
(8.1) $\big\{\text{localizing $\otimes$-ideals of $\mathscr{T}$}\big\}$$\big\{\text{subsets of $\operatorname{Spc}^{h}(\mathscr{T}^{c})$}\big\}$$\big\{\text{subsets of $\operatorname{Spc}(\mathscr{T}^{c})$}\big\}$$\scriptstyle{\operatorname{Supp}^{h}}$$\scriptstyle{\operatorname{Supp}}$$\scriptstyle{\pi}$
@@
For each $i\in I$, the $t$-category $\mathscr{S}_{i}$ is tt-stratified by assumption, hence h-stratified by 8.6. Descent for h-stratification as established in 7.22 then implies that $\mathscr{T}$ is h-stratified. Therefore, 8.6 implies that $\mathscr{T}$ is tt-stratified, due to our assumption that $\mathscr{T}$ satisfies Nerves of Steel. ∎
@@
The equivalence between $(a)$ and $(b)$ is 8.7. 7.12 gives $(b)\Rightarrow(c)$, while the reverse implication is 7.22. ∎
@@ 9. Descent for the Nerves of Steel conjecture
where the vertical arrows are the finite localizations and the bottom arrow is the induced functor. This diagram satisfies the Beck–Chevalley condition $h^{*}f_{*}\simeq g_{*}k^{*}$; cf. [SAN22, Lemma 5.5]. Indeed, this may be checked after applying the conservative $h_{*}$. For any $s\in\mathscr{S}$, we have $h_{*}h^{*}f_{*}(s)\simeq{f_{Y}\otimes f_{*}(s)}\simeq{f_{*}(f^{*}(f_{Y})\otimes s)}\simeq f_{*}(f_{\varphi^{-1}(Y)}\otimes s)\simeq f_{*}(k_{*}k^{*}(\mathbb{1})\otimes s)\simeq f_{*}k_{*}k^{*}(s)\simeq h_{*}g_{*}k^{*}(s)$ where we have invoked [BS17, Proposition 5.11]. With this in hand, consider a compact object $x\in\mathscr{S}(\varphi^{-1}(V))^{c}$. Then $x\oplus\Sigma x=k^{*}(c)$ for some $c\in\mathscr{S}^{c}$ and it suffices to show that $g_{*}k^{*}(c)$ is compact. This follows from our hypothesis, since $g_{*}k^{*}(c)\simeq h^{*}f_{*}(c)$. ∎
@@
$\bigsqcup_{i\in I}\operatorname{Spc}^{h}(\mathscr{S}_{i}^{c})$$\operatorname{Spc}^{h}(\mathscr{T}^{c})$$\bigsqcup_{i\in I}\operatorname{Spc}(\mathscr{S}_{i}^{c})$$\operatorname{Spc}(\mathscr{T}^{c}).$$\scriptstyle{\sqcup\pi_{\mathscr{S}_{i}}}\scriptstyle{\simeq}$$\scriptstyle{\varphi^{h}}$$\scriptstyle{\pi_{\mathscr{T}}}$$\scriptstyle{\varphi}$
Recall from 7.9 that $(f_{i}^{*})$ being jointly nil-conservative implies that $\varphi$ is also surjective. If $\varphi$ is in addition injective as assumed in $(a)$, then it is bijective. Since $\sqcup\pi_{\mathscr{S}_{i}}$ is bijective, $\varphi^{h}$ is injective and hence bijective. Therefore $\pi_{\mathscr{T}}$ is bijective.
Now assume either condition $(b)$ or $(c)$. By 6.14, the Nerves of Steel conjecture holds also for any finite localization of $\mathscr{S}_{i}$. Moreover, the property of being strongly closed is local in the target by 9.1, while the property of being nil-conservative is local in the target by 7.9[$(c)\Leftrightarrow(d)$] and 6.15. Therefore, we can reduce to the case when $\mathscr{T}$ is local and check that the fiber of $\pi_{\mathscr{T}}\colon\operatorname{Spc}^{h}(\mathscr{T}^{c})\to\operatorname{Spc}(\mathscr{T}^{c})$ over the unique closed point $\mathfrak{m}\in\operatorname{Spc}(\mathscr{T}^{c})$ is a singleton. Since $\varphi$ is surjective, we may choose $i\in I$ and a closed point $\mathscr{M}\in\operatorname{Spc}(\mathscr{S}_{i}^{c})$ lying over $\mathfrak{m}$. For the remainder of this proof, we set $\mathscr{S}=\mathscr{S}_{i}$ and simplify the notation accordingly.
@@
and hence $f^{*}(E_{\mathscr{B}})\otimes z=0$ because homological support detects weak rings (1.8). Therefore, $E_{\mathscr{B}}\otimes f_{*}(z)=0$ by the projection formula. That is $\mathscr{B}\not\in\operatorname{Supp}^{h}(f_{*}(z))$. By the hypothesis that $f$ is strongly closed, $f_{*}(z)$ is compact and hence $\operatorname{Supp}^{h}(f_{*}(z))=\pi_{\mathscr{T}}^{-1}(\operatorname{supp}(f_{*}(z)))$ by 5.2. Therefore we have $\mathscr{B}\not\in\pi_{\mathscr{T}}^{-1}(\operatorname{supp}(f_{*}(z)))$ so $\mathfrak{m}=\pi_{\mathscr{T}}(\mathscr{B})\not\in\operatorname{supp}(f_{*}(z))=\varphi(\operatorname{supp}(z))$ where the last equality is by [BCH+23b, Theorem 13.13] since $z$ is a compact weak ring. But $\varphi(\operatorname{supp}(z))$ contains $\varphi(\{\mathscr{M}\})=\mathfrak{m}$ yielding the desired contradiction. ∎
@@
The results of 9.2 are not entirely satisfying; variations on the theme are possible and we expect a more complete statement to exist. Nevertheless, using 9.2(a)(a), we can provide a variant of the argument used in [BAL20b] to deduce the Nerves of Steel conjecture from nilpotence-type theorems in several examples of interest:
@@
Let $\mathscr{T}$ be a rigidly-compactly generated tt-category. Suppose there exists a tt-field $f_{\mathscr{P}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{P}}$ detecting $\mathscr{P}$ for each prime $\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})$, i.e., $\mathscr{P}$ is in the image of $\operatorname{Spc}(f_{\mathscr{P}}^{*})$. The following are equivalent:
@@
$\mathscr{T}$$\text{Mod-}\mathscr{T}^{c}=\mathscr{A}$$\mkern 1.5mu\overline{\mkern-1.5mu\mathscr{A}\mkern-1.5mu}\mkern 1.5mu_{\mathscr{B}}$$\mathscr{F}_{\mathscr{P}}$$\text{Mod-}\mathscr{F}_{\mathscr{P}}^{c}$$\scriptstyle{\overline{\mathbb{h}}_{\mathscr{B}}}$$\scriptstyle{f_{\mathscr{P}}^{*}}$$\scriptstyle{\mathbb{h}}$$\scriptstyle{\mkern 1.5mu\overline{\mkern-1.5muF\mkern-1.5mu}\mkern 1.5mu}$$\scriptstyle{\mathbb{h}}$
@@
$\bigsqcup_{\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})}\operatorname{Spc}^{h}(\mathscr{F}_{\mathscr{P}}^{c})$$\operatorname{Spc}^{h}(\mathscr{T}^{c})$$\bigsqcup_{\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})}\operatorname{Spc}(\mathscr{F}_{\mathscr{P}}^{c})$$\operatorname{Spc}(\mathscr{T}^{c}),$$\scriptstyle{\sqcup\pi_{\mathscr{F}_{\mathscr{P}}}}\scriptstyle{\simeq}$$\scriptstyle{\varphi^{h}}$$\scriptstyle{\pi_{\mathscr{T}}}$$\scriptstyle{\varphi}\scriptstyle{\simeq}$
we see that $(a)$ implies that the map $\varphi^{h}\colon\bigsqcup_{\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})}\operatorname{Spc}^{h}(\mathscr{F}_{\mathscr{P}}^{c})\to\operatorname{Spc}^{h}(\mathscr{T}^{c})$ is surjective. This is equivalent to the family being nil-conservative by [BCH+24b, Theorem 1.9]. Thus $(a)\Rightarrow(c)$. Finally, the implication $(c)\Rightarrow(a)$ follows from 9.2(a). ∎
@@
The next result establishes an equivalent formulation for the Nerves of Steel conjecture entirely in terms of the tensor triangular support. It strengthens previous work [BHS23c, Proposition 3.13] and could also be used to give an alternative deduction of 8.6.
@@
Assume first that the Nerves of Steel conjecture holds for $\mathscr{T}$. Using 5.11 twice as well as the tensor product formula for homological support (1.1), we compute
@@
Conversely, suppose that the tensor product formula holds for all weak rings. The argument in the proof of [BHS23c, Proposition 3.13] then implies the Nerves of Steel conjecture for $\mathscr{T}$. Note that the result there assumed the full tensor product formula, but it is only used for the weak rings $E_{\mathscr{B}}$. ∎
@@ 10. Applications to tt-stratification
We now state our main applications. We start with three descent results which extend those of [BCH+24b, Section 17] and [BCH+24a, Section 12] to a family of functors.
@@
By 8.6 each $\mathscr{S}_{i}$ is h-stratified and satisfies the Nerves of Steel conjecture. The result then follows from 9.2 and 8.8 since $(f_{i}^{*})_{i\in I}$ is jointly nil-conservative by 7.9. ∎
@@
In the situation of 10.1(c)(c), if the indexing set $I$ is finite then $\operatorname{Spc}(\mathscr{T}^{c})$ is automatically noetherian and hence the hypothesis that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian is superfluous.
@@
Since $A$ is descendable, the base-change functor $A\otimes-\colon\mathscr{C}\to\operatorname{Mod}_{\mathscr{C}}(A)$ is conservative and hence induces a surjection $\operatorname{Spc}(\operatorname{Mod}_{\mathscr{C}}(A)^{c})\to\operatorname{Spc}(\mathscr{C}^{c})$; see [BCH+24b, Theorem 1.4] or [BCH+24a, Proposition 7.8]. Hence $\operatorname{Spc}(\mathscr{C}^{c})$ is also noetherian. We may then invoke 10.1(c)(c). ∎
@@
The previous corollary strengthens the results of [BCH+24a, Section 12] by removing the hypothesis that the descendable algebra $A$ be separable of finite tt-degree.
@@
$\bigsqcup_{i\in I}\operatorname{Spc}^{h}(\mathscr{S}_{i}^{c})$$\operatorname{Spc}^{h}(\mathscr{T}^{c})$$\bigsqcup_{i\in I}\operatorname{Spc}(\mathscr{S}_{i}^{c})$$\operatorname{Spc}(\mathscr{T}^{c})$$\bigsqcup_{i\in I}\operatorname{Spec}^{h}(\operatorname{End}^{\bullet}_{\mathscr{S}_{i}}(\mathbb{1}))$$\operatorname{Spec}^{h}(\operatorname{End}^{\bullet}_{\mathscr{T}}(\mathbb{1}))$$\scriptstyle{\sqcup\pi_{\mathscr{S}_{i}}}$$\scriptstyle{\pi_{\mathscr{T}}}$$\scriptstyle{\sqcup\rho_{\mathscr{S}_{i}}}$$\scriptstyle{\rho_{\mathscr{T}}}\scriptstyle{\simeq}$
in which the three indicated vertical maps are surjections. The top and middle horizontal maps are also surjections by 7.9. The bottom map is a bijection by assumption, hence $\rho_{\mathscr{T}}$ is surjective as well. Since the tt-categories $\mathscr{S}_{i}$ are cohomologically stratified, the left vertical composite is a bijection, which then implies that all maps in this diagram are bijections. The result thus follows from 8.8. ∎